OpenStudy (anonymous):

6 e^x = 3 e^x + 15 Find the exact value of x. No decimal entries.

OpenStudy (anonymous):

What's ^ stand for?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

my bad I meant the power I was reading something about roots

OpenStudy (anonymous):

x=1−ln(2) x≈0.30685281944

OpenStudy (anonymous):

lets first divide each side by 3. now we have 2e^x = e^x + 5. Bringing over the e^x to the other side we get e^x = 5. Now using natural logs:= lne^x = x = ln 5

OpenStudy (anonymous):

got it instead of writing ln I was writing log thanks